OpenStudy (anonymous):

e^x=2 how do you solve for x?

OpenStudy (anonymous):

write it in logarithmic form.... \(\large b^x=y \leftrightarrow x=log_by \)

OpenStudy (anonymous):

in this case use natural log

OpenStudy (anonymous):

Apply natural logarithm in both sides of the equation \[\ln(e^x)=\ln2\] By the logarithm properties, \[xln(e)=\ln2\] and lne=1, so finally x=ln2